# On the Peril of (Even a Little) Nonstationarity in Satisficing Regret Minimization

### 19 Mar 2026 v1

@@ 6.4 Proof of Theorem 4
In this section, we prove Theorem 4 by adapting the proof of Theorem 1 in Michel et al. (2022) from the stationary setting to our nonstationary setting under Assumptions 2 and 3. We work with the same regret representation as in equation (19): $R_{T}^{S}(\pi;\nu)=\sum_{\ell=0}^{L-1}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a).$ Following the event decomposition used in Appendix A of Michel et al. (2022), we write, for each arm $a$ and time $t$, $\left\{a_{t}=a\right\}\subset\{t=a\}\cup\left\{a_{t}=a,Z_{t}^{c}\right\}\cup\left\{a_{t}=a,Z_{t}\right\},$ where $Z_{t}:=\left\{\forall a\in[K],\hat{\mu}_{t}(a)<S\right\}$. For the first event, we have
$\sum_{\ell=1}^{L}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(t=a)\leq\sum_{a:\Delta_{a,\min}>0}\Delta_{a,\max}.$
For each arm $a\in[K]\setminus a^{*}$, define $T(a):=\max\{t\in[T]:\mu_{t}(a)<S\}$. For $n\leq T(a)$, let $\hat{\mu}_{a,n}$ denote the empirical mean of arm $a$ based on the first $n$ pulls of $a$ that occur strictly before time $T(a)$. For the second event, Assumption 3 implies that
$\sum_{\ell=1}^{L}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a,Z_{t}^{c})$ $\leq\sum_{a:\Delta_{a,\min}>0}\Delta_{a,\max}\sum_{n=1}^{T(a)}\mathbb{P}(\hat{\mu}_{a,n}(a)\geq S)$
$\leq\sum_{a:\Delta_{a,\min}>0}\Delta_{a,\max}\sum_{n=1}^{T(a)}\exp(-\frac{n\Delta_{a,\min}^{2}}{2})\lesssim\sum_{a:\Delta_{a,\min}>0}\frac{\Delta_{a,\max}}{\Delta_{a,\min}^{2}},$
Define $\hat{\mu}_{a^{*},n}$ denote the empirical mean of arm $a^{*}$ based on the first $n$ pulls of $a^{*}$. For the last event, Assumption 2 implies that
$\sum_{\ell=1}^{L}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a,Z_{t})$ $=\sum_{\ell=1}^{L}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a^{*},Z_{t})$
$\leq\sum_{a:\Delta_{a,\min}>0}\Delta_{a,\max}\sum_{n=1}^{T}\mathbb{P}(\hat{\mu}_{a^{*},n}\leq S)\lesssim\sum_{a:\Delta_{a,\min}>0}\frac{\Delta_{a,\max}}{\Delta_{*}^{2}}.$
Combining three events together, we complete the proof of Theorem 4.
@@ 7 Conclusion
This paper studies satisficing regret in nonstationary bandits. Our results reveal a sharp contrast with the stationary realizable setting: once the reward process is allowed to change at unknown times, constant satisficing regret is no longer achievable in general. In particular, in the realizable piecewise-stationary setting with $L$ stationary segments, we show that the optimal regret scales as $\Theta(L\log T)$ whenever $L\geq 2$.
Our lower bounds are based on a new post-interaction-reference framework, which strictly extends interactive Fano methods from stationary bandits to nonstationary settings.
We also show that constant satisficing regret can still be recovered under additional structural assumptions. An important direction for future work is to characterize more precisely which structural conditions are sufficient, and which are necessary, for constant satisficing regret under nonstationarity.
@@
Q. Xie and Y. Zhang are supported in part by National Science Foundation (NSF) grants CNS-1955997 and EPCN-2339794 and EPCN-2432546.

### 25 Apr 2026 v2

@@ 6.4 Proof of Theorem 4
In this section, we prove Theorem 4 by adapting the proof of Theorem 1 in Michel et al. (2022) from the stationary setting to our nonstationary setting under Assumptions 2 and 3. We work with the same regret representation as in equation (19): $R_{T}^{S}(\pi;\nu)=\sum_{\ell=0}^{L-1}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a).$ Following the event decomposition used in Appendix A of Michel et al. (2022), we write, for each arm $a$ and time $t$, $\left\{a_{t}=a\right\}\subset\{t=a\}\cup\left\{a_{t}=a,Z_{t}^{c}\right\}\cup\left\{a_{t}=a,Z_{t}\right\},$ where $Z_{t}:=\left\{\forall a\in[K],\hat{\mu}_{t}(a)<S\right\}$. For the first event, we have $\sum_{\ell=0}^{L-1}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathds{1}\{t=a\}\leq\sum_{a:\Delta_{a,\min}>0}\Delta_{a,\max}.$
For each arm $a\in[K]\setminus\{a^{*}\}$, define $T(a):=\max\{t\in[T]:\mu_{t}(a)<S\}$. For $n\leq T(a)$, let $\hat{\mu}_{a,n}$ denote the empirical mean of arm $a$ based on the first $n$ pulls of $a$ that occur strictly before time $T(a)$. For the second event, Assumption 3 implies that
$\sum_{\ell=0}^{L-1}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a,Z_{t}^{c})$ $\leq\sum_{a:\Delta_{a,\min}>0}\Delta_{a,\max}\sum_{n=1}^{T(a)}\mathbb{P}(\hat{\mu}_{a,n}(a)\geq S)$
$\leq\sum_{a:\Delta_{a,\min}>0}\Delta_{a,\max}\sum_{n=1}^{T(a)}\exp(-\frac{n\Delta_{a,\min}^{2}}{2})\lesssim\sum_{a:\Delta_{a,\min}>0}\frac{\Delta_{a,\max}}{\Delta_{a,\min}^{2}}.$
Let $\hat{\mu}_{a^{*},n}$ denote the empirical mean of arm $a^{*}$ based on the first $n$ pulls of $a^{*}$. For the last event, a similar argument, together with Assumption 2, implies that
$\sum_{\ell=0}^{L-1}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a,Z_{t})$ $=\sum_{\ell=1}^{L}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a^{*},Z_{t})\lesssim\sum_{a:\Delta_{a,\min}>0}\frac{\Delta_{a,\max}}{\Delta_{*}^{2}}.$
Combining the three events, we complete the proof of Theorem 4.
@@ 7 Conclusion
This paper studies satisficing regret in nonstationary bandits. Our results reveal a sharp contrast with the stationary realizable setting: under nonstationarity, even a single change in the environment is enough to make $T$-independent satisficing regret unattainable in general. In particular, in the realizable piecewise-stationary setting with $L$ stationary segments, we show that the optimal regret scales as $\Theta(L\log T)$ whenever $L\geq 2$. Our lower bounds are based on a new post-interaction reference framework, together with a conditional Fano argument, that strictly extends the interactive Fano methods from stationary bandits to the nonstationary setting.